OpenStudy (anonymous):

What point in the feasible region maximizes P for the objective function P = 2x + 3y? Constraints: 2x+y≤15 x+3y≤20 x≥0, y≥0 ive tried for so long but i have no idea.. answer choices: A. (0,15) B. (5,5) C. (8,7) D. (2,6)

jimthompson5910 (jim_thompson5910):

are you able to graph the inequalities?

OpenStudy (anonymous):

yes, I did that but i think i might of done it wrong..

OpenStudy (amistre64):

can you draw it out?

OpenStudy (anonymous):



OpenStudy (anonymous):

this is what I got. I'm not sure if it's right becasue arent the intersection points supposed to be the answers?

jimthompson5910 (jim_thompson5910):

that's a good rough sketch, so you have the right graph where does each line intersect? In other words, what are the vertices of the boundaries?

OpenStudy (amistre64):

the intersections are the testng points yes

OpenStudy (anonymous):

the points from the intersections were not apart of the answer choices though , so i don't know what to do

OpenStudy (amistre64):

now, if the options given are not a point of intersection, then you simply want to choose the one that is in the boundaries and makes P maximal

OpenStudy (anonymous):

all of the answer choices are like 6.6666 for x or 5.76 or some not whole number and in the answer choices, all the numbers are whole

OpenStudy (amistre64):

we can tell if a point is in the feasible region if it meets all of the constraints

OpenStudy (amistre64):

for example: 0,15 is fine for eq1, but fails for eq2 ... its not in the region

OpenStudy (anonymous):

so it doesn't have to be the exact intersecting point? it can just be in the intersection region?

OpenStudy (amistre64):

its asking which option is in the feasible region, and maximizes P

OpenStudy (amistre64):

first step, determine which options are in the feasible region, then test them out for maximum

OpenStudy (anonymous):

Ohh okay, thank you!!

OpenStudy (amistre64):

A and C do not meet the criterias, so we need to test the remaining points

OpenStudy (anonymous):

I think it's b??

OpenStudy (amistre64):

10+15 is bigger then 4+18 yeah

OpenStudy (anonymous):

okay great, thanks so much

OpenStudy (amistre64):

youre welcome

OpenStudy (phi):

Here is the graph (I used Geogebra, which you can download for free) Evaluate P at each "intersection point" point (5,5) maximizes P= 2x+3y= 10+15= 25